How does y=(-1/2)x+3 change into x+2y-6=0?
Answer by stanbon(75887)

How does y=(-1/2)x+3 change into x+2y-6=0?
-----
Multiply both sides by 2 to get:
2y = -x + 6
-----
Rearrange to get:
x + 2y - 6 = 0
==================
Cheers,
Stan H.
==================